OpenStudy (anonymous):

Please help! I need help with substitution method! 1. x+y=8 y=3x 2. 7x-2y=1 2y=x-1 I don't need to know how to solve it, I already know, I just have lots of work to do and my deadline is in like 2 hours! Thanks for your help and understanding :)

OpenStudy (anonymous):

@UsukiDoll, you really seem to like Rarity from MLP! ;P

OpenStudy (usukidoll):

subsitute y = 3x in x+y =8

OpenStudy (anonymous):

I know that, I just need the steps

OpenStudy (usukidoll):

then add x + 3x = 8 4x = 8 x = 2

OpenStudy (usukidoll):

subsitute 2 in the y= 3 x DATS 6

OpenStudy (usukidoll):

so ok ya x = 2 and y = 6 for the first 1

OpenStudy (anonymous):

Thx! :)

OpenStudy (usukidoll):

7x-2y=1 2y=x-1 now umm letz seeeeee - x from both sides

OpenStudy (usukidoll):

make it all preettyyy yahhhh

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Do u watch MLP???

OpenStudy (usukidoll):

7x-2y=1 2y-x = -1 Rearrangeee a bit yahhhh -x+2y = -1 7x -2y = 1 oooo datz nice we can cancel the y;sss

OpenStudy (usukidoll):

yeah

OpenStudy (usukidoll):

6x = 0 datz ah 0 so y = 1/2 0-2y = 1 divide by -2 y = -1/2

OpenStudy (anonymous):

So lemme guess, Rarity is ur fav Pony? ;P

OpenStudy (usukidoll):

-x+2y = -1 0 + 2(-1/2) = -1 -1 = -1 7x -2y = 1 0 - 2 (-1/2) = 1 1 = 1 so yahhh dey werk n twerk

OpenStudy (usukidoll):

umm yeah I got the plush and 3 in hasbro figure

OpenStudy (anonymous):

Nice!

OpenStudy (anonymous):

SO what's x and y?

OpenStudy (usukidoll):

y = -1/2 x = 0 for second probz

OpenStudy (anonymous):

Thx so much! Do u know the elimination method?

OpenStudy (usukidoll):

ya... but subsitution ez bettah ... matrix method es mi favorita.

OpenStudy (anonymous):

Ya but I'm supposed to solve this one using elimination method... 3x+3y=27 x-3y=-11

OpenStudy (usukidoll):

efff de instructions same result appears

OpenStudy (anonymous):

And 2x+y=4 6x+7y=12 using substitution

OpenStudy (usukidoll):

ooo in dat case get rid of the y

OpenStudy (usukidoll):

3x+3y=27 x-3y=-11 y kan cancelz

OpenStudy (usukidoll):

then u get 4x = 16

OpenStudy (anonymous):

Then???

OpenStudy (usukidoll):

x = 4 and then sub it back to one equation and get y

OpenStudy (anonymous):

So y is....

OpenStudy (usukidoll):

I eating at da moment

OpenStudy (anonymous):

Sorry if it seems like I'm pushing the answers out of ya it's just that I wanna finish this and get rid of it ASAP

OpenStudy (anonymous):

Oh sorry then...

OpenStudy (usukidoll):

ah I got 3x+3y=27 diz es 12 + 3y = 27 subtraxt u get 15 3y= 15 Y= 5

OpenStudy (usukidoll):

y = 5 x = 444444

OpenStudy (usukidoll):

I mean 4

OpenStudy (anonymous):

Thx! :D

OpenStudy (anonymous):

Sorry 4 da trubble

OpenStudy (anonymous):

(Trouble)

OpenStudy (usukidoll):

2x+y=4 6x+7y=12 using substitution ya u mutiply -3 for the first equation

OpenStudy (usukidoll):

so ummmmmmmmmmmm -6x-3y=-12 6x + 7y = 12 easy way to get de y 4y = 0 y = 0

OpenStudy (usukidoll):

sub bak in hea 6x + 7y = 12 6x = 12 x = 2 so x = 2 y = 0 and give me medal for this

OpenStudy (anonymous):

So how do I give medal???

OpenStudy (anonymous):

I'm new so ya

OpenStudy (usukidoll):

click on best response button

OpenStudy (anonymous):

Did it

OpenStudy (anonymous):

THX